UT 20170924 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (LBTO)
Observer AO-Team: DMiller, GTaylor (LBTO)
Telescope Operator: Sallenson
LBTO Support: TSargent

Observations done remotely from Tucson

Plan for the night

Same as last night:

- AO-tests

- LUCI-AO short images

- LUCI-AO AFC if time is left

Summary

- Lots of AO-tests, many issues

- Excercising LUCI-AO short images not really good AO-wise (lots of issues)

- AO system seems to be very fragile

Details

All times are UT
Files are luci1/2.20170924.NNNN.fits

01:20 Sunset, sky is clear, T = +4deg, hum = 60%, wind ~ 10 m/s

Looks like it will be a good night

AO-group is setting up their systems

JHill is putting in a new collimation model

02:14 Steve checks pointing and collimates, seeing looks moderat (about 1.2" - 1.5")

02:25 AO-tests/troubleshooting starts, they do tests in Bin = 1 and 2

02:40 DIMM wanders between 1.3 and 1.8"

Taking LUCI2 night-sky spectra in parallel

N3.75, LS1.5_2.0,

# 2: Throughslit, 10sec
# 3: K, 21502, 60sec
# 4: H, 16727, 60sec
# 5: J, 12902, 60sec
# 6: z, 9791, 60sec

N30, N30LS _0.13arcsec with 2 holes

# 7: Throughslit, 10sec
# 8: K, 21502, 300sec
off by 200pix, aligned mirror
# 9: H, 16727, 300sec
redo alignment
# 10: H, 16727, 300sec
adjusted alignment in spectral direction (-48AA)
# 11: H, 16727, 300sec
adjusted alignment in spectral direction once more (-9AA)
# 12: H, 16727, 300sec
# 13: H, 16727, 300sec, redo with srr, 11 samples, better
# 14: K, 21502, 300sec, srr,
# 15: J, 12902, 300sec, srr
# 16: H, 16727, 300sec, srr, brightest lines easily detected (S/N 13 + 6)

N30, N30LS _0.13arcsec with 25 holes

SRR; 300sec only

Lines and S/N for lines are from BA Umstaetter

# 17: H, 16727, brightest lines (S/N 13 + 6) easily detected
# 18: H, 16549, S/N 2.5 detected
# 19: H, 16409, S/N 1.5 detected

DIMM reports constantly 1" - 1.5", some clouds with extinction of about 1.5mag passing

# 20: Dark
# 21: J, 1.2348, S/N 1.5 detected, S/N 0.5 probably not, seems to be sort of threshold (TBC)
# 22: J, 1.2118

Serious problems with the AO@SX:
-> the shell can hardly be set
-> for whatever reason the gain has to be limited, otherwise the system won't work

07:30 Clouds are getting thicker, humidity rises (over 80% now)

AO-tests are done, we are close to the humidity limit (above 90%), try to
start LUCI-AO bino scripts.

Short images, offsets within Patrol field

Bin=2

08:38 Preset to 3C 452, H-band, OK, RunAO, OK, start taking data

SX use 400 Hz, DX uses 625 Hz

DIMM reports below 1", clouds have gone, AO-weather !

08:51 Loop opened @ SX, (was luckily paused), closed manually, cont.,

09:13 Loop opened @ SX, image L1#35 crap, cont'd

09:20 Loop opened @ DX, star not found on CCD39/47 after offset, manually recovered, cont.
same failure @ DX again,

09:31 Resent preset, failed @ SX (interrupted system call), LUCI still thinks preset is running and waits until
the DX-side has closed the loop and than returns an error message,

09:42 Resent preset, OK, RunAO, OK,

09:50 Offset @ SX failed (same as for DX @ 09:20), manually recovered, OK

09:53 Safe skip @ SX (DIMM = 1.0"), man. recovered, cont'd,

10:00 Loop opened @ SX, stopped, target too low

L1: # 21 - 38 L2: # 34 - 50

Bin = 2

10:03 Preset to SU Aur, Fe II and Br_gam and H2, failed on DX (large slew), Steve corrects pointing,
resent preset, OK, RunAO, failed due to seeing...waiting a bit and close loop

NCPA on @ SX, NCPA mostly off @ DX

Jump in seeing wrt to previous position, DIMM = 1.4 -> 2.8 -> 0.9" !!!

10:42 Loop open @ DX (NOT due to seeing), recovered manually, cont'd

10:55 Loop open @ DX (NOT due to seeing), recovered manually, cont'd

10:58 Offset failed @ DX (arbitrator issue, NOT due to seeing),

On some of the images SU Aur looks like a binary due to reflexes/ghosts

L1: # 40 - 74
L2: # 52 - 86

Bin = 1

11:27 Preset to Trapezium, OK, RunAO, SX OK, shell RIPed @ DX, RunAO hangs somewhere @ DX,
restart arbitrator, loop opened @ SX as there was a seeing-fluke

11:49 Resent preset, OK, RunAO, seeing spike up to 1.5", stopped

Handover to Dave for an ESM-test at the very bitter end of night

-- %USERSIG{JochenHeidt - 2017-09-24}%
Topic revision: r7 - 24 Sep 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback